Instrument Log

25M MFR / US / UIR            C1 CF Lamont, OK

Date

Technician

Comments

6/20/16

JEM

Maintenance was being performed on the 25M T/H from 1715 GMT to 1730 GMT. Data may be effected during this time.  RTS @1730 GMT.

6/14/16 – 6/15/16

CAW/TMG

SE tower booms needed to be lowered to install new cameras on tower. Lowered SE booms at 1800 GMT on 6/14.  Booms remained down overnight while working on camera mounts.  Raised booms after cameras were installed and RTS at 2025 GMT on 6/15.

12/3/15

JEM

DQPR #4987 requests that OS v24 and Logger Program v1.4 be loaded.  Lowered the Southwest tower booms @1524 GMT. Updated the OS to version 24 @1535 GMT. Loaded program version 1.4 @1543 GMT. Raised the Southwest booms back into place @1602 and RTS @1605 GMT.  US and UIR data will also be affected during this time.

11/3/15

JEM

A new program needs to be loaded onto the data logger to fix an issue with Logger ID's larger than 7999 not showing up properly.  Loaded the new program @1820 GMT.  Verified readings were all correct.  RTS @1825 GMT.

10/27/15

JEM

The UIR and US need to be replaced for calibration.  Lowered the SE boom @1522 GMT.  R/R the *UIR and *US.  Edited and reloaded the datalogger program @1600 GMT.  Created a gap in the data from 1602 GMT to 1612 GMT to clearly show when the change out occurred.  Raised the SE boom and RTS @1619 GMT.

*Old SNs:   UIR– 30682   US – 30793

*New SNs:  UIR – 30341  US  – 30939

10/26/15

CAM

Lowered the SE booms at 1435 GMT for T/H calibration changes.  Raised booms and RTS at 1949 GMT.  No problem with this instrument.

8/25/15

JEM

In DQPR #4782 the mentor indicates that the 25 MFR heater board failed on 20150729 and requests its replacement.  Lowered the SE booms at 1835 GMT.  R/R the heater board and raised the booms and RTS @1850 GMT.  Data will also be affected for the UIR and US during this time that the booms were lowered.

8/26/15

CAW

25M MFR due 90 day calibration checks.  Lowered tower at 1220 GMT and ran tower back up at 1230 GMT while calibrating MFR head MP-248.  Lowered tower at 1940 GMT to reinstall the MFR head and ran tower back up at 1950.

4/9/15

PD

PGS shed at the base of the 60M tower is being replaced with a 20' ISO container.  Crane and other equipment used to remove the PGS shed may have been in the field of vision of this instrument between 1420 and 1610 GMT.  All equipment was out of the field of vision by 1610 GMT.

2/2/15

JEM

The 25M MFR needs to be reinstalled.  Lowered the boom @ 2048 GMT.  Reinstalled the sensor head and raised the boom.  RTS @2100 GMT.

1/30/15

JEM

Since the previous maintenance on the 25M MFR, all diffusers are reporting -9999.  Lowered the boom @1932 GMT.  Found the 0.5A diffuser fuse was blown.  Replaced fuse and verified readings.  Went ahead and removed the sensor head so that Craig could do his 90-day calibration on the sensor.  Sensor will be reinstalled on 2/2/15.

1/27/15

JEM

According to DQPR #4442, the MFR25M 415 nm wavelength is dropping out.  Lowered the boom @2106 GMT and reseated all logger connections as well as all sensor cable connections. Also checked the sensor head connector for corrosion, but none was found.  Raised the booms @2119 GMT.  RTS @2120 GMT.

10/28/14

CAM

Lowered the SE booms at 1520 GMT for T/H calibration changes.  Raised booms and RTS at 2230 GMT.  No problem with this instrument.

8/27/14

JEM

The 25M US needs to be replaced for annual calibration.  Lowered the East boom @1330 GMT.  R/R the *US and loaded the new program @1341 GMT.  Verified serial number and cal factor.  Raised the East boom back into place @1348 GMT.  RTS @1350 GMT.

*Old - 30802     *New – 30793

6/23/14

CAM

No data had been collected from 25MFR since 1209 GMT on 6/21.  Found the UPS that powers the comms equipment in the AOS trailer was down.  Power cycled UPS and restored power at 1220 GMT.  Normal data collections resumed at next interval. 

6/5/14

CAM

25 Meter MFR due 90 day calibration check.  Brought tower down and removed MFR head for 90 day calibration. Down at 1635 and back up at 1645 on 6/4/14.  Completed calibration and reinstalled MFR head. tower down at 1420 and back up at 1435 on 6/5/14.  System returned to service.

5/27/14

CAM

No data had been collected from instrument since 1209 GMT on 5/25.  Found the UPS that powers the comms equipment in the AOS trailer was down.  Breaker for UPS supply line was tripped.  Reset breaker at 1325 GMT.  UPS powered up and comms equipment power restored.  Normal data collections resumed at next interval. 

2/25/14

CAW

25 Meter MFR due 90 calibration check.  Lowered tower at 1440 GMT.
Remover MFR head and raised tower to operational position at 1500 GMT,  Ran MFR head through calibration checks, lowered tower at 2045 GMT and reinstalled MFR head,  Raised tower to operational position at 2100 GMT.

6/3/13

CAW

25 meter MFR due 90 day calibration check.  Lowered tower & removed MFR head for calibration at 1410 GM on 6/3.  Found several bad channels so replaced old MHR head with new head and installed 6\6\13 at 1625 GMT.  *Old – 915  *New - 248

3/4/13

CAW

25 meter MFR due 90 day calibration check.  Tower lowered at 1430 GMT and MFR head removed for 90 day calibration check.  MFR head reinstalled and system returned to service at 2030 GMT.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Power was restored at 1920 GMT.  Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/8/13

JEM

DQPR #3616 requests the 25M IRT lens be replaced.  Lowered the SE booms @1942 GMT.  R/R the IRT lens @1948 GMT.  Raised the SE booms and RTS @1957 GMT.

12/24/12

JEM

Tower booms need to be lowered to clean the 25M IRT len and verify IRT settings.  Tower was lowered @1544 GMT.  Raised the tower @1627 GMT.  Status OK.  RTS @1630 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  Power was restored at 2115 GMT after electrical upgrades were complete. 

11/13/2012

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Tower instruments resumed normal operation.

10/31/12

JEM

The boom needs to be lowered to verify the 25M IRT serial number.  Booms were lowered form 1513 GMT to 1535 GMT.

11/5/12

CAW

Lowered booms at 1338 GMT for 25 MFR 90-day calibration.  Raised booms and RTS at 1349 GMT after MFR head removed.  Lowered booms again for reinstallation at 2025 GMT.  Raised booms and RTS at 2036 GMT.

10/9/12

CAM

Lowered the SE booms at 1900 GMT for T/H calibration changes.  Raised booms and RTS at 2145 GMT.  No problem with this instrument.

7/26/12

CAW

25 Meter MFR Head due for 90 day calibration check.  Lowered tower at 1405 GMT and removed MFR head.  Did 90 day calibration check.  Reinstalled MFR head and RTS at 2000 GMT.

7/13/2012

JEM

The 25M *US is scheduled to be replaced for calibration.  Lowered the SE tower elevator @1943 GMT.  Disconnected the US @1947 GMT.  Replaced the sensor and reconnected @1956 GMT.  Raised the SE tower elevator @2002 GMT.  Loaded the new program @2021 GMT.  Verified readings and RTS @2023 GMT.

*Old -  30673F3  *New - 30947F3

6/26/12

CAM/PAD

SDS is not collecting data from 25M IRT.  Powercycled side A 25M boom of tower at 1726 GMT with no change in status.  Lowered side A booms at 1837 GMT.  Raised booms at 2007 GMT.  No issues know with instrument at this time.  RTS at 2007 GMT.

6/4/12

JEM

25M MFR lowered so 60M T/RH troubleshooting can be performed.  Lowered the SE booms at 1410 GMT.   Raised the SE booms @1637 GMT and RTS @1640 GMT.

6/1/12

JEM

25M MFR lowered so 60M T/RH sensor can be replaced.  Lowered the SE booms at 1412 GMT.   Raised the SE booms and RTS @1512 GMT.

5/24/12

PAD

DQPR 3407 states there is a low bias for 60 T/H probe on side A of tower.  Mentor suspects bad connection and suggests lowering/raising tower.  Lowered side A 25 and 60m booms at 2115 GMT.  All indicator lights for signal and AC were lit after raising booms.  No issues known with instrument at this time.  RTS at 2130 GMT.

5/8/12

CAW

25M MFR is due for 90-day calibratoin.  Lowered booms at 1258 GMT and removed MFR.  Raised booms and RTS at 1310 GMT.  Lowered booms again for reinstallation at 1522 GMT.  Raised booms and RTS at 1535 GMT.

3/8/12

PAD

Radon Gas Analyzer inlet pipe has separated approximately 40M above the ground.  Contractor is on site today to climb tower and repair inlet tube.  Power to tower needs to be shut down before contractor can climb.  Shut down power at base of tower to 25 and 60M levels, sides A and B at 2023 GMT.  After repairs were made and complete inspection of inlet pipe was performed, power was restored.  No issues with this instrument are known at this time.  RTS at 2257 GMT.

2/24/12

JEM

The Mentor has provided a new program for the 25M MFR.  Loaded the new program @1724 GMT.  Readings observed normal.  RTS @1745 GMT.

2/8/12

JEM

The *25M MFR is scheduled for an upgrade.  Lowered the tower @1855 GMT.  Removed the old control box and components, MFR head and mounting hardware.  Installed new mounting hardware, MFR head, enclosure and components.  Loaded the program and tested @2155 GMT.  Readings appeared normal.  The tower was raised back into place at 2212 GMT.  Communications were tested and readings were normal at 2215 GMT.   SDS will have to update collections to be able to collect from a CR1000. 

*Old SN’s:  Head - 248  Logger - BE67  Control Box - WD27206

*New SN’s:  Head - 370  CR1000 Logger - 1766  CR1000 WP - 1910  Power Supply - RB06107424

The new control box does not have a WD sticker.

11/1/11

CAM

Lowered the SE booms at 1815 GMT for T/H calibration changes.  Raised booms and RTS at 2130 GMT.  No problem with this instrument.

7/19/11

CAW

25M *US needs to be changed out with replacement instrument.  Lowered tower at 1910 GMT,   25M US was installed, new program was loaded on data logger and tower was raised at 2107 GMT.  RTS at 2110 GMT.   *Old - 30939F3  *New - 30673F3

7/19/11

CAW

25M MFRSR is due for 90 day calibration.  Lowered tower at 1910 GMT and removed instrument for calibration.  Installed instrument and raised tower at 2107 GMT.  RTS at 2110 GMT.

4/15/11

JEM

The instrument has not collected since the 0400 hour.  Due to other instruments that are down the source of the problems seems to be the Rocketport hub.  The Rocketport power light is flashing.  Power cycling the Rocketport hub did not correct the issue.  Notified the IT department of the issue.

4/7/11

CAW

MFR is due for calibration.  Lowered tower at 1200 GMT and raised tower at 1215 GMT.  Calibrated instrument.  Lowered tower at 1433 GMT, installed instrument, and raised tower at 1444 GMT.

11/23/10

CAW

MFR is due for 90-day calibration.  Removed MFR head at 1825 GMT and calibrated.  Reinstalled and RTS at 2010 GMT.

10/19/10

CAM

Tower Systems on site for tower inspections.  Need to secure power for maintenance.  Secured AC power to 25m and 60m levels at 1407 GMT.  Restored power following inspection at 1620 GMT.

10/18/10

CAM

Lowered the SE booms at 1445 GMT for T/H calibration changes.  Raised booms and RTS at 2013 GMT.  No problem with this instrument.

8/19/2010

JEM

The 25M US needs to be replaced for calibration.  Lowered the southeast boom @1430 GMT.  R/R the *US @1440 GMT.  Loaded the new program to the datalogger @1443 GMT.  Raised the southeast boom back into position @1451 GMT.  RTS @1500 GMT.

*Old – 31097F3     *New – 30939F3    

8/19/2010

JEM

The 25MFR head is ready to be reinstalled.  Lowered the southeast boom @1430 GMT.  Installed the MFR head @1435 GMT.  Raised the southeast boom back into position @1451 GMT.  RTS @1500 GMT.

8/18/2010

JEM

The 25MFR head needs to be removed for calibration.  Lowered the southeast boom @1826 GMT.  Removed the 25MFR head for calibration.  Raised the southeast boom back into position @1826 GMT.  RTS @1841 GMT.

7/14/10

CAM/JEM

60M tower SE booms need to be lowered to remove GPS array repeater.  Lowered SE booms at 1445 GMT.  Removed GPS equipment.  Raised booms and RTS at 1556 GMT.

5/11/10

CAW

25M MFR was removed at 1848 GMT for calibration and needs to be installed.  Lowered side A booms at 2040 GMT and installed MFR.  Booms raised and system RTS at 2055 GMT.

5/11/10

CAW

25M MFR is due for calibration.  Lowered tower, removed MFR head at 1848 GMT and calibrated.  Raised tower at 1900 GMT.

2/16/10

CAM

25M UIR and 25M US desiccant needed to be replaced.  Replaced desiccant while boom was down to remove MFR for calibration.  RTS at 1507 GMT.

2/16/10

CAM

25M MFR needs to be removed for calibration.  Lowered SE booms at 1455 GMT and removed MFR.  Raised booms at 1507 GMT.  Lowered booms again at 1712 GMT to reinstall MFR after it was calibrated.  Raised booms and RTS at 1722 GMT.

11/19/09

CAM

SDS instrument status page indicated no data had been collected since 2109 GMT on 11/18.  T/S found the UPS in the AOS trailer which powers the tower communications equipment had no output.  AC LINE and ALARM LEDs were lit on the UPS.  Power cycled UPS at 1342 GMT and power was restored.  RTS at 1345 GMT. 

11/9/09

CAM

25M MFR needs to be removed for calibration.  Lowered SE booms at 1405 GMT and removed MFR.  Raised booms at 1414 GMT.  Lowered booms again at 1623 GMT to reinstall MFR after it was calibrated.  Raised booms and RTS at 1633 GMT.

9/21/09

CAM

Lowered the SE booms at 1450 GMT for T/H calibration changes.  Raised booms and RTS at 1735 GMT.  No problem with this instrument.

8/26/09

CAW

25M MFR needs to be reinstalled.  Lowered booms and installed MFR at 1930 GMT.

8/25/09

CAW

25M MFR head had no response when placed on the spectral bench.  Need to remove the head for troubleshooting.  Lowered SE booms at 1320 GMT and removed head.  Raised booms at 1330 GMT.

8/19/09

MK

The 25M MFR needs to be removed and calibrated. Lowered tower on 8/19/09 at 1725-1738 GMT removed MFR head and started calibrations. On 8/20/09 calibrations were completed, lowered tower at 1455 GMT-1510 GMT reinstalled MFR.

5/19/09

CAM

The ingest has been failing since the MFR was reinstalled after calibration on 5/13.  The log file shows the head ID is reporting as 0.   Lowered SE booms at 1743 GMT.  Checked MFR head cable and pins.  No problem found.  Reseated cable, raised booms and RTS at 1758 GMT.  Check of status later determined problem still exists.  Will likely need to replace sensor next.

5/13/09       

CAM

25M MFR needs to be reinstalled after it was calibrated.  Lowered booms at 2146 GMT and installed MFR.  Replaced US and UIR desiccant at 2152 GMT.  Raised booms and RTS at 2201 GMT.

5/13/09

CAM

25M MFR needs to be removed for calibration.  Lowered booms at 1431 GMT and removed MFR.  Raised booms and RTS at 1446 GMT.

2/12/09

MK

25m MFR was removed for calibration and needs to be installed.  Lowered tower, side A, at 1610 GMT.  Installed instrument and raised tower.  RTS at 1640 GMT.

2/4/09

MK

Mentor has requested lamp calibrations for the 25M MFR.  Lowered tower at 1656 GMT. Removed 25M MFR, raised tower at 1707 GMT. The MFRSR head has been taken to the calibration facility so Craig Webb can calibrate on 2/5/09.

10/1/08

CAM

SE booms were lowered at 1850 GMT to clean IRT.  Observed the 25M US and 25M UIR desiccant needed to be changed.  Replaced desiccant in US and UIR at 1855 GMT.  Raised booms and RTS at 1905 GMT.

9/12/08

JEM

The US needs to be replaced for annual calibration.  Lowered the tower @1455 GMT.  R/R the *US radiometer.  Raised the tower @1520 GMT.  The program was not loaded until 1825 GMT.  Verified S/N's and cal factors.  RTS @1830 GMT.   *Old - 30939F3    *New - 30945F3 

12/4/07

MK

The 25M MFR is due for a standard lamp calibration. Lowered tower 1950 GMT removed MFR for standard lamp calibration. Raised tower at 2050 GMT.

7/26/07

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1333 GMT.  Replaced desiccant in the UIR while booms were down.  Also replaced the *US radiometer for annual calibration.  Raised booms and RTS at 1920 GMT.                               *Old – 30652F3  *New – 30939F3

2/5/07

MK

The 25M MFR is scheduled for lamp calibration. Lowered tower at 2025 GMT, ran lamp calibrations, reinstalled at raised tower at 2115  GMT.

10/11/06

CAM

The 25M US radiometer is due for calibration.  SE booms were lowered at 1447 GMT for 25m/60m T/H calibrations.  R/R the *US radiometer while the booms were down.  Downloaded the new program into the data logger.  Raised booms and RTS at 2141 GMT.    

*Old - 30953F3     *New - 30652F3.

10/11/06

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1447 GMT.  Raised booms and RTS at 2141 GMT.

5/31/06

PAD

Jane McKinney request 60m tower lowered to record model and serial numbers from the wireless data transceiver located on SE 60M boom.  Lowered side A of tower at 1755 GMT.  Raised tower at 1851 GMT.  No problem with this instrument.  RTS at 1855 GMT. 

3/27/06

PAD

SDS has not collected data from 25M IRT since 3/25/06 at 1403 GMT.  No problems with this instrument at this time.  Power cycled 25M instruments on side A of tower at 1615 GMT.  RTS at 1630 GMT.

8/5/05

MR

The 25M US is due for calibration.  R/R the *US.  Downloaded the new program into the data logger.  RTS at 1600 GMT.  Cal factors=8.98 10-6      *Old - 30948    *New - 30953.

8/5/05

MR

The 25M UIR is due for calibration. R/R the *UIR.  Downloaded the new program into the data logger.  RTS at 1600 GMT.  Cal Factors k0=-14.01, k1=.24181, k2=1.027, k3=-2.378  *Old – 30835F3   *New – 30682F3**

8/5/05

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1327 GMT.  Raised booms and RTS at 1637 GMT.

8/4/05

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Tower Systems techs also on site for tower inspection.  Turned the 25 and 60M levels power off at 1330 GMT so Tower Systems could service tower. Lowered booms at 1830 GMT.  Raised booms and RTS at 2014 GMT.

6/10/05

MR

The SE tower needs to be lowered due to maintenance on other equipment.  Lowered the tower at 1945 GMT.  Raised the tower at 2045 GMT.  RTS at 2100 GMT.

9/27/04

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1438 GMT.  Raised booms and RTS at 2056 GMT.

8/5/04

MR

The mentor wants the MFR sensor head replaced. Lowered the tower at 1410 GMT.  R/R the *MFR sensor head.  Raised the tower at 1445 GMT.  RTS at 1450 GMT.  *Old - 917   *New - 248

8/5/04

MR

The UIR sensor needs to be replaced. Lowered the tower at 1410 GMT.  R/R the *UIR sensor.  Raised the tower at 1445 GMT.  RTS at 1450.   *Old - 30784F3  New - 30835F3

7/21/04

MR

The tower South east needs to be lowered and the instruments need to be cleaned due to the farmer working the ground under the tower.  Lowered the elevator at 1526 GMT.  Cleaned the sensor at 1605 GMT.  Raised the elevator at 1615 GMT.  RTS at 1626 GMT.

7/6/04

PAD

25 MFR last successful collection was on 7/2/04 at 2302 GMT.  Power cycled 25M instruments on side A (S.E. side) of tower at 1419 GMT.  Instrument successfully collected at 1602 GMT.  RTS at 1434 GMT.

4/16/04

CAM

Need to power cycle the SE 25M level power due to IRT communications problem.  There is no failure with this instrument.  Power cycled the SE 25M level power at 2035 GMT. 

4/15/04

CM\PD

Observed while the booms were lowered for MFR maintenance that the US and UIR desiccant needed to be replaced.  R/R US and UIR desiccant.  Booms were lowered at 2104 GMT and raised at 2151 GMT.

4/15/04

CM/PD

DQPR  267 states the 25M MFR showed small peaks in the data on March 6-8 around 1300 GMT and again on March 31 and April 1 around 2300 GMT.  DQPR suggests possibly something like a cable in the view of the sensor.  Lowered the SE booms at 2104 GMT for inspection.  Observed nothing directly beneath the sensor but the US radiometer’s data cable was hanging low enough to be in the 180 degree plane of the MFR.  Secured the US cable above the MFR.  Raised SE booms and RTS at 2151 GMT.

2/23/04

MR

The SE tower connections need to be reseated for other equipment.  Lowered the elevator about 1 foot and the raised the elevator.  RTS at 1853 GMT.

2/20/04

MR

The 25M MFR needs to be reinstalled as per mentors instructions.  He will send a new one out as soon as possible but wants the old one reinstalled even though the channels are noisy.  Lowered the tower at 1445 GMT.  Reinstalled the MFR head and data logger.  Raised the tower at 1515 GMT.  Notified Ron Reed and he was able to communicate and force the logger to start ok.  RTS at 1530 GMT.

2/13/04

MR

The instrument mentor wants the MFR head taken off the tower and put on the LICOR calibrator.  The channels are reading on the high side.  Lowered the elevator at 1752.  Removed the MFR head and data logger and put them on the LICOR calibrator.  The readings were OK except for channel 2 was erratic.  Notified the mentor and he had me remove the connector on the back and check for corrosion.  Found a little corrosion on the connector and cleaned it off.  Put the head back on the calibrator but this did not correct the high reading problem.  He said to leave the instrument off the elevator for now and he will check into sending a new head.  Raised the elevator at 2134.  RTS at 2140 GMT.

2/10/04

MR

The elevator needs to be lowered due to maintenance on other equipment.  Lowered the elevator at 1905 GMT.   Raised the elevator at 2315 GMT.  RTS at 2330 GMT.

1/15/04

MR

The program needs to be changed and uploaded so that it reflects the new UIR calibration coefficient changes. Uploaded the new program into the data logger at 2052 GMT.  RTS at 2105 GMT.

12/8/03

MR

The tower needs to be lowered and the calibration needs to be run on the 25M MFR.  Lowered the tower at 1940. Performed the calibration on the head and board.  Raised the tower at 2105.  RTS at 2105.

11/19/03

MR

The 25M and 60 M tower elevator needs to be lowered due to maintenance on other equipment.  Lowered the SE tower at 2105.  Raised the tower at 2120 GMT.

10/7/03

MR

The tower connections need to be reset.   Lowered the tower about 3 inches at 1530.  Raised the tower at 1531.  RTS at 1545.

10/6/03

MR

The tower needs to be lowered due to maintenance on other instruments. Lowered the tower at 1713 GMT.  Raised the tower at 1826.  RTS at 1828 GMT.

8/19/03

MR

The mentor requests that the MFR data logger and head be removed and a LICOR calibration be run on them.  Lowered the tower at 1415 GMT.  Removed the head and data logger.  Performed a calibration on the head and data logger.  Reinstalled the head and logger.  Sent the results to the mentor.  Raised the tower at 1605.  RTS at 1615.

8/15/03

CAM

Need to lower SE booms for annual 25 and 60M T/H changeouts.  Lowered the SE booms at 1341 GMT for maintenance.  Raised booms and RTS at 1920 GMT.

8/14/03

CM

SE booms lowered a few inches at 1900 GMT to tension the cable between the ground and 25M boom.  Raised booms at 1914 GMT.  Only the 60M boom made connection at that time due to the added tension.  Lowered booms again at 1925 GMT and repositioned the 25M boom higher on the cable.  Raised booms and RTS both 25 and 60M levels at 1935 GMT.

8/14/03

CM

SE booms lowered for tower system maintenance.  There was no failure with this instrument.  Lowered boom at 1443 GMT.  Raised boom and RTS at 1515 GMT.

8/5/03

MR

The elevator tower needs to be lowered due to maintenance on other equipment.  Lowered the tower at 1836 GMT.  Raised the tower at 1910 GMT.  RTS at 1921 GMT.

5/16/03

JEM

Instruments on 25M East boom were power cycled @2047 GMT due to problem with 25M MFR.  No problems with this instrument.  RTS @ 2100 GMT.

5/16/03

JEM

The 25M MFR has not been sending files.  Power cycled the 25M MFR @2047 GMT.  RTS @2100 GMT.

3/21/03

MR

The readings on the US are reading a negative value.  Lowered the tower at 1824.  According to the pin out of the US plug pin A should be wired to 4L channel on the data logger and pin B should be wired to 4H on the data logger.  Testing the pin out with a multimeter it was discovered that pin A and B where switched at the data logger end of the cable and this was causing the negative readings.  Changed the 4H and 4L wires on the data logger and the US reading returned to a positive reading.  RTS at 2024 GMT.

3/21/03

MR

The instrument mentor for the IRT requests that the wiring and program be checked for the IRT/UIR/US due to the dome temp not matching the IRT temp.  T/S found that the shielding wire for the UIR was not connected through the multiplexer.  Connected the shield wire to the correct channel.  Called Vic Morris and consulted with him what channels I needed to check to see if the reading for the temperature was correct.  He requested that channels 5, 9 and 10 should read about the same.  Channels 5, 9 and 10 where reading 298.46, 294.21 and 294.00 respectively.  Notified Vic Morris and he said that should be about right and wanted the tower raised back up and he would monitor the channels to see if the  shield wire made a difference.  Raised the tower at 2024 GMT.

3/18/03

JEM

The 25M MFR, US and UIR have not collected since 3/17/03 @1502 GMT.  The power strip that powers the bit driver in the AOS trailer was found powered off.  Powered the power strip back on @1505 GMT.  The power strip was likely powered off by accident while rearranging the AOS system.  RTS @1515 GMT.

3/17/03

MR

The instrument mentor wants the IRT and US/UIR connections checked. Checked the connections.  They checked good.  Due to approaching thunderstorm we had to raise the tower.  Will continue with this CM when weather permits.  Raised the tower at 1705 GMT.

2/21/03

MR

Due to maintenance on other equipment the elevator needs to be lowered.  Lowered the elevator at 1953 GMT.  Raised the elevator at 2056 GMT.  RTS at 2056 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

1/30/03

JEM

The 60M Southeast temperature has been too high.  There is no problem with the 25M MFR.  Corrective maintenance is being done on the 60M T/H.  Lowered Southeast 25M and 60M booms @1655 GMT.  RTS @1810 GMT.

12/16/02

MR

The serial numbers and calibration factors in the program for the UIR and US need to be checked and changed if necessary.  Lowered the tower at 1639 GMT.  The program was retrieved, checked and the calibration factors where changed.  Raised the tower at 1705 GMT.  RTS at 1710.

11/7/02

CM/MR

SE booms lowered from 1557 GMT to 1755 GMT to troubleshoot 25M IRT communications problems.

11/7/02

CM/MR

SDS comlog reports indicated that the 25MFR collections had been failing since the booms were lowered on 11/6.  T/S found a failed bit driver on the rocket port hub in the AOS trailer.  R/R bit driver and communications RTS at 1540 GMT. 

11/6/02

CM/MR

SE booms lowered from 2017 GMT to 2222 GMT to troubleshoot 25M IRT communications problems.

10/31/02

MR

The instrument mentor wants the 25 M MFR checked.  All the channels have started reading bad.  T/S confirmed that the channels where bad.  The power supply checks ok and the board voltage checked ok.  R/R the *data logger.  The channel readings returned to normal.  RTS at 1616 GMT.  *Old - 93959  *New - BE67

10/29/02

CAM

SE booms lowered from 2135 GMT to 2220 GMT for 25 and 60M T/H maintenance.

10/28/02

CAM

The SE 25 and 60M T/H sensors are reading –9999’s or 0’s.  Lowered the SE booms slightly at 1731 GMT and then raised to reseat connections.  RTS at 1733 GMT.

10/25/02

MR

The instrument mentor needs the MFRSR data logger and head put back on.    Lowered the tower at 1805 GMT.   Put the original head and data logger back in the system.  The head cable is suspected as being bad.  Mentor will send a new cable to be installed.  RTS at 1920 GMT.

10/24/02

MR

The instrument mentor wants the 25 M MFR data logger and head removed for testing.  Lowered the tower at 1415 GMT.  Removed the *data logger and **head.  Raised the tower at 1430 GMT.  data logger *Old - 93959  **New - 917

9/23/02

MR

The instrument mentor wants the head connections checked for corrosion due to the heater voltage randomly dropping out.  Lowered the tower at 1930 GMT.  Checked the head cable at the head and saw no corrosion in the either connector.  Replaced the connector and checked the connector on the board and it checked good.  Raised the tower at 1943 GMT.  RTS at 1945 GMT.

9/10/02

MR

The instrument mentor wants the tower lowered so work can be done on the PGS sampling system.  Lowered the tower at 1733 GMT.  Raised tower at 1837 GMT.  Status ok.  RTS at 1848.

9/3/02

MR

The 25M elevator needs to be lowered to check an instrument for the PGS system.    Lowered the tower at 1940 GMT.  Raised the tower at 2030.  RTS at 2040.

8/2/02

D. Cook

Booms lowered from 1354 to 1912 GMT for tower maintenance and 25/60m T/H calibrations. 

8/1/02

MR

The *PSP needs to be changed out due to the calibrations being overdue.  Removed the old PSP and installed the new PSP.  Changed the calibration coefficients in the program.  RTS at 1645 GMT.  *Old - 30671F3 CF=116.2   *New - 30948F3 CF=117.79

8/1/02

MR

The *PIR needs to be changed out due to the calibrations being overdue.  Removed the old PIR and installed the new PIR.  Changed the calibration coefficients in the program.  RTS at 1645 GMT.   *Old - 29665F3 CF=251.89   *New - 30784F3 CF=267.38 

8/1/02

D. Cook

TSI on site.  Booms lowered from 1300 to 2130 GMT for tower maintenance. 

5/3/02

CAM

According to the quicklook data plots, the 25M MFR is once again not reporting head temperature or solar data.  Time of outage appears to be around 0400 GMT on 5/2.   60M tower operations are suspended at this time due to faults with the tower.  No corrective action taken at this time.

4/11/02

CAM/MR

Instrument mentor advised that the 25M MFR data still looked bad.  Lowered the 25M MFR at 2003 GMT.  Connected a laptop directly to the data logger and monitored the data and it looked good.  Checked all signal cable connections at the 25M level and at the tower base to be sure they were all tight.  Raised the 25M MFR and monitored from the base of the tower.  The data looked good from there as well.  RTS at 2106 GMT.  Checked the 25M MFR quicklook plots the following day and the data now appears to be good.   Suspect a signal connection was originally not making good contact and reseating/tightening the connections corrected the problem.

4/2/02

CAM

MFR mentor reports there has been no solar data from the 25 MFR since 3/24 and no head temperature data sine 3/25.  The SE booms needed to be lowered momentarily to reseat the MFR connections.    Lowered the SE booms at 1754 GMT.  Raised the booms and RTS at 1755 GMT.

2/19/02

MR

The 25M MFRSR has not sent any data to the SDS since the power outage on 1/30/02.  T/S found that the instrument was working but had not been started.  Further T/S found that both bit drivers in the communications line where locked up.  Power cycled both bit drivers at 1604 GMT.  RTS at 1615 GMT.

2/13/02

MR

The tower needs to be lowered so the MFRR can be power cycled.  Lowered the tower at 1500.  Raised the tower at 1505.  RTS at 1515 GMT.

12/19/01

MR

The tower needs to be lowered due to calibration on the PGS system. Lowered the tower at 2055 GMT.  Raised the tower at 2110 GMT.

11/01/01

MR

Vic Morris requested that the upwelling wiring be checked due to the readings on the US and UIR are bad. Lowered the tower at 2050 GMT.  T/S found that the 100K precision resistor between the CR10 and multiplexer was not installed.  Installed the resistor and notified the mentor of the change.  Raised the tower at 2103 GMT.

10/4/01

MR

Joe Michalsky wants to run the calibrations on the 25M MFR a second time.  Lowered the east side of the tower at 1828 GMT.  Raised the tower at 1900.  RTS at 1905 GMT.

10/3/01

MR

Joe Michalsky is on site and needs to perform a calibration on the 25M MFR.  Lowered the tower at 2006.  Performed the calibration on the MFR with the LICOR calibrator.  Raised the tower at 2057.  RTS at 2105 GMT.

10/2/01

MR

The instrument mentor notified us that the 25 meter MFR is not sending files.  Lowered the 25 meter tower.  The mentor also notified us that the interface board can be removed.  Removed the board.  Raised the tower at 1600.  RTS at 1603 GMT.

9/27/01

MR

The IRT, UIR, US is being separated from the MFRSR.  Lowered the 25 and 60 meter tower at 1503.  Removed the UIR, IRT and US from the MFRSR board and installed them on their own data logger.  Raised the tower at 1912.

8/15/01

CAW

25M US radiometer needs changed out with a newly calibrated one.  Lowered SE booms at 1530 GMT.  R/R *US radiometer and RTS at 1555 GMT.   *Old - 30810F3   *New - 30671F3

8/7/01-8/8/01

MR

Southeast booms lowered at 1918 GMT on 8/7 for 25M and 60M T/H calibrations.  Booms were left down at ground level overnight.  Raised booms and RTS at 1840 GMT on 8/8.

8/7/01

MR

Tower systems is on site to perform maintenance on the 25M elevator and the 60M tower. Lowered the south east 25M elevator.  Tower systems replaced one of the connectors on the tower and performed general maintenance on the elevator and tower components.  RTS at 1910 GMT.

7/20/01

MR

The tower will be lowered due to work being performed on the PGS system.  No work will be performed on this instrument. Lowered tower at 1450 GMT.  Performed PGS system maintenance.  Raised tower at 1700 GMT.

7/13/01

MR

The 25m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2000 GMT.  Marc Fisher performed maintenance on the PGS system.  Raised the tower at 2130.  RTS at 2145 GMT.

7/13/01

MR

The SDS is unable to talk to the 25M T/H and 60M T/H southeast. Lowered the tower to power cycle the instruments at 1440 GMT.  Raised the tower at 1441 GMT.  Communications were restored.

7/12/01

MR

The 25m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2100 GMT.  Marc Fisher performed maintenance on the PGS system.  Raised the tower at 2135.  RTS at 2145 GMT.

5/18/01

MR

The instrument mentor requests that the data logger be checked for the mux chips due to the instrument reporting 00's for the IRT and upwelling radiometers. Lowered the tower at 1707 GMT. Checked for the MUX chips and observed they where missing. Installed the chips. Raised the tower and RTS at 1719 GMT.

5/4/01

MR

The data logger needs to be replaced for the system upgrade. Lowered the tower elevator at 1839 GMT.  R/R the *data logger.  Raised the tower and RTS at 1854 GMT.  *Old - 93957 *New - 93959

11/27/00

CAW

Lowered the SE 25m and 60m booms from 1840 to 2035 GMT to clean the chilled mirrors.

9/21/2000

CAM

David Cook requested that the SE 25 and 60 meter booms be lowered and reseated as part of troubleshooting bad humidity sensor readings. Lowered the booms slightly at 1850 GMT. Raised the booms and RTS at 1852 GMT.

1/6/2000

CM/CW

25M and 60M booms lowered from 1941 GMT to 2123 GMT for installation of new GPS antenna. Cleaned radiometer domes while the booms were down.

11/16/99

CM/SS

The A side booms were lowered from 1700 GMT to 1820 GMT to finish insulating the 25M IRT.

11/15/99

CM/GS

The A side booms were lowered from 2011 GMT to 2217 GMT to insulate the 25M IRT.

7/2/99

ZW

The 25M US was due for change out for calibration. At 1720 Site Ops lowered the 25M platform. R/R the 25M *US. Raised the 25M platform. Status ok. RTS 1750 GMT. *Old - 30901F3 *New - 30958F3

5/19/99

5/20/99

CAM/DC

The 25M boom was lowered at 1755 GMT on 5/19/99 for T/H maintenance. This resulted in the MFR, US, and UIR being lowered as well. The boom was left with sensors connected at ground level overnight. During that time it was noted that the 25M US and UIR desiccant needed to be replaced. Replaced 25M US and UIR desiccant at 1600 GMT on 5/20/99. Raised boom and RTS at 1915 GMT.

12/8/98

CM/GS

The 25m boom (side A) was lowered at 2105 GMT for IRT maintenance. This resulted in the 25MFR, US, and UIR being taken off line as well. Raised the booms and RTS at 2153 GMT.

9/25/98

ZW

The 25M US was due to be calibrated. Lowered the 25M tower at 1700 GMT. R/R the 25M *US. Raised the tower. Status ok. RTS 1730 GMT. *Old - 31634F3 *New - 30901F3

8/11/98

ZW

About 1/2 of a day after the 25M MFR Data Logger was installed it stopped communication with the SDS system. At 1850 GMT Powered off the 25M MFR, and R/R the *Data Logger. Powered 25M MFR on at 1900 GMT. Status ok. RTS 1910 GMT. *Old - 93909 *New - 93957

8/6/98

GRS

Lowered the 60m tower booms on side A at 1920 GMT to read some program values on the 25m IRT. Raised the booms and RTS the instrument at 1939 GMT.

8/3/98

ZW

The MFR Mentor came to calibrate the 25M MFR sensor, and Data Logger board. At 1945 GMT Site Ops lowered the tower and R/R the 25M *MFR sensor head. Also R/R the **Data Logger. Powered on system and raised the tower at 2045 GMT. Status ok. RTS 2045 GMT. *Old - 991 *New - 917c **Old - 93957 **New - 93909

7/16/98

7/17/98

CAM/DC

All tower booms were lowered from 1458 GMT on 7/16/98 until 1513 GMT on 7/17/98 for 25M and 60M T/H calibrations. The booms were powered at ground level during that time.

6/23/98

MW

SDS operations reported a loss of communications with all of the 60M tower instruments beginning at 1004 GMT today. Power cycled the tower COS at 1404 GMT. Communications restored after the power cycle. RTS at 1405 GMT.

5/18/98

CAM

SDS operator advised that the 25 MFR collections had failed on the last three attempts. The last successful collection was 1435 GMT on 5/18. The SDS comslog indicates 'Could not get data from logger'. T/S found that the 60 meter tower breaker was tripped at the tower base. Reset breaker but it immediately tripped again. Turned all instrument levels AC power off and reset breaker again. The breaker stayed on this time. Powered the instruments AC power on one level at a time and isolated the fault to the 60 meter A side. Lowered the A side booms at 1945 GMT. Found that the AC power connections were loose inside the junction box at the 60 meter level and had arced together thus tripping the breaker. Reconnected the AC connections and tightened the lugs. Connected short AC power jumpers at ground level and tested. AC power status now O.K. Raised booms at 2035 GMT. Attempted to link with all instruments through the COS but was still unable to link with the 25 MFR. Power cycled MFR at 2105 GMT. Was then able to receive a hello back from the data logger. RTS status O.K. at 2115 GMT.

5/4/98

CAM

SDS TR indicated that no files had been received from the 25M MFR since 1813 GMT on Friday, May 1. The error message 'Login prompt not received' was reported. Tipped the COS but was unable to receive 'Hello' prompt from data logger. Placed a loop back on the bit driver at the tower and verified that the fibers and bit drivers were O.K. Power cycled the 25M level power. Still no response from data logger. Lowered the booms slightly and then raised to reseat the connections at 1530 GMT. Was then able to receive a 'Hello' back from data logger at 1540 GMT. RTS status O.K.

5/1/98

CM/MW

The A side 25M and 60M booms were lowered at 1826 GMT for chilled mirror removal. Raised booms and RTS at 1935 GMT.

2/6/98

MW

SDS ops indicated that no files were being collected from the 25M MFR since 2011 on 2/4/98. Reset the GFCI at 2122 GMT. RTS.

1/28/98

MW

SDS ops reported a loss of communications with the 25M MFR starting at 1811 GMT. Found the GFCI was tripped. Reset GFCI at 2217 GMT. RTS.

1/21/98

CAM

SDS TR indicated no files had been received from 25M MFR since 2011 GMT on 1/20/98. T/S found that the GFCI breaker was tripped at the base of the tower for the 'A' lower side. Reset breaker at 1525 GMT on 1/21/98. SDS reinitialized data logger at 1611 GMT and collections resumed status O.K. at 1711 GMT on 1/21/98.

1/14/98

MW

The 25M MFR GFCI was tripped. The SDS has not been getting data from the 25M MFR since 2210 GMT on 1/11/98. Reset the GFCI at 1710 GMT. Data collection resumed at 1711 GMT.

1/7/98

CAM

SDS TR indicated that no files had been received from the 25M MFR since 0935 GMT on 1/4/98. T/S on 1/7/98 found that the 'A' side 25M GFCI breaker at the base of the tower was tripped. Reset GFCI at 2035 GMT on 1/7. SDS collection resumed at 2131 GMT. RTS status O.K.

12/19/97

CAM

Lowered booms from 2100 GMT to 2200 GMT for chilled mirror maintenance.

11/19/97

CAM

SDS TR indicated no files being received from 25 MFR since 0113 GMT today. T/S found that the GFCI circuit breaker at the base of the tower was tripped. Reset breaker at 1718 GMT. Was then able to receive a hello back from data logger through the COS. SDS reinitialized the 25 MFR at 1813 GMT. RTS status O.K.

11/13/97

MW

SDS operations reported a loss of communications with the 25M MFR instrument. The 25 M GFCI was tripped. Reset the GFCI at 1728 GMT. RTS.

10/15/97

MW

The SDS stopped collecting files from the 25M MFR starting at 0409 GMT on 10/12/97. The message "Login prompt not received" was displayed in the COMLOG. Reset the 25M GFCI at 1340 GMT. SDS initialized the logger at 1416 GMT. RTS.

9/27/97

DKB

IOP reports no files from the 25 MFR, US or UIR. Reset GFCI breaker at 0245 GMT.

9/26/97

CAM

South booms lowered from 2010 to 2047 GMT to R/R 25M T/H.

9/24/97

MW

SDS operations reported a loss of communications with the 25M MFR at 0508 GMT. Found the 25M GFCI was tripped. Reset the GFCI at 1806 GMT. Power restored to the 25M instruments.

9/12/97

CAW

25M US/PSP out of calibration and needs exchanged. *R/R US/PSP. * Old – 29280 * New - 31634

9/10/97

DN

Lowered the tower from 1915 to 1950 GMT for IOP instrument installation preliminary.

9/04/97

CAM

Scott Richardson was on site and needed to lower the east (A) booms of the tower for planning and preparation of mounting instruments on the booms for the upcoming IOP. Lowered the east (A) 25 and 60 meter booms at 1415 GMT for measurements. Raised booms and RTS at 1445 GMT.

8/14/97

GRS

Raised the 25 boom at 1549 GMT. RTS the MFR, US, UIR.

8/13/97

GRS

The 25m MFR sensor head was loosened from its mount for calibration purposes. At 1715 GMT, the 25 MFR sensor was removed from its normal mount (still plugged in) for calibration. The sensor head was cleaned from 1715-1716 GMT. Placed the sensor in the optical radiation calibrator at 1716 GMT. Switched the calibrator on at 1719 GMT. Performed the required checks and calibrations. Re-installed the sensor 1730 GMT. RTS at 1730 GMT. NOTE: The 25 MFR was already at ground level due to other installs on the 25m boom.

8/12/97

GRS

Lowered the boom at 1625 GMT for temp/hum install and calibrations. Collections are off and the instrument will remain down until calibrations are complete.

10/21/96

CAM

Lowered boom for training from 1740 to 1800 GMT.

10/18/96

CAM

Lowered boom to re-seat UIR connections to correct noisy signal reported by mentor. RTS at 1900 GMT.

10/8/96

CAM/GRS

Lowered tower and * R/R MFR data logger with 5 MHz board in continuing T/S effort to correct short data files. RTS at 2003 GMT. * old - 93909 * new - 93957

9/20/96

CAW/JS

Lowered boom to R/R MFR E-Prom at 1420 GMT.

9/10/96

CAM/JS

Lowered boom to install IOP instruments at 1440 GMT. R/R US and UIR desiccant at this time.

4/22/96

CAM/ZW

1810 GMT Lowered tower to T/S 25 IRT. RTS at 1900 GMT.

4/19/96

CAW/CAM

1330 GMT Lowered tower and disconnected 25 IRT at mentors request. RTS at 1340 GMT.

4/18/96

CAW/CAM

1412 GMT Lowered tower to T/S 25 IRT. Found loose connection and corrected. RTS at 1433 GMT.

4/8/96

SDS DMT

T/S found "ICM" file missing on the SDS. Replaced and tested OK at 2110 GMT.

4/7/96

SDS DMT

Discovered corrupted directory structure. Corrected but still no files being received by SDS. T/S continues.

4/5/96

CAM

1600 GMT T/S no files received by SDS since 0332 GMT. Reset bit drivers and SDS was able to TIP the MFR logger at 1605 GMT. Still no files being received by SDS. Lowered tower and re-seated all connections and power cycled system at 2110 GMT. T/S will continue.

3/7/96

TMG/CAM

2005 GMT. Lowered tower to * install E-Prom upgrade. R/R UIR and US desiccant. RTS at 2031 GMT. * old - Rev J * new - Rev L

3/1/96

CAM

T/S continues. Lowered sensor and checked everything. DMT informed us that they had stopped collection to perform testing. Collection started up and tested OK. RTS at 1731 GMT.

2/29/96

CAM

1800 GMT. T/S why SDS has not received files since 2/28 2330 GMT. Lowered sensor and raised to reset the tower connection with no luck. SDS can link with logger but no files being collected.

2/13/96

CAM/GRS

AC power secured for construction at 1510 GMT. RTS at 2149 GMT.

1/29/96

TMG/CAM

AC power secured for construction at 1515 GMT. RTS at 2030 GMT.

 12/27/95

TMG/CAM

AC power secured for construction at 1530 GMT.